You are choosing between two long distance telephone plans. Plan A has a monthly fee of $15 with a charge of $0.08 per minute for all long distance calls. Plan B has a monthly fee of $3.00 with a charge of $0.12 per minute for long distance calls. For how many minutes of long distance will the cost for the two plans be the same and what will be the cost of the plans? If make approximately 15 long distance calls per month averaging 30 minutes per call which plan should you select? explain?
Answer by mananth(16949)

You are choosing between two long distance telephone plans. Plan A has a monthly fee of $15 with a charge of $0.08 per minute for all long distance calls. Plan B has a monthly fee of $3.00 with a charge of $0.12 per minute for long distance calls. For how many minutes of long distance will the cost for the two plans be the same and what will be the cost of the plans? If make approximately 15 long distance calls per month averaging 30 minutes per call which plan should you select? explain?
...
Let the time in minutes be t per month
Plan A 15+0.08t
Plan B 3+0.12t
..
PLANS TO BE EQUAL
15+0.08t=3+0.12t
0.12t-0.08t=15-3
0.04t=12
/0.04
t= 12/0.04
t= 300 minutes they will be equal..
..
Plan A 15+0.08t
Plan B 3+0.12t
15 calls 30 minutes each = 450 minutes
..
Plan A 15+0.08t
15+0.08*450
=15+36
=$49
Plan B 3+0.12t
3+0.12*450
3+54=57
=$57
..
You know which to choose.
...
m.ananth@hotmail.ca